Calcium carbonate (Caco,) which is found in limestone and marble reacts readily with hydrochloric acid (HCI) to produce calcium chloride (CaCl). carbon dioxide, and water.


a. Write the balanced chemical reaction:


b. What mass of hydrochloric acid will be needed to produce 325 g of

calcium chloride?


c. What mass of carbon dioxide could be produced if 750 g of calcium carbonate reacts with excess hydrochloric acid?



1
Expert's answer
2021-05-27T07:34:25-0400

a. CaCO3 + 2HCl = CaCl2 + H2O + CO2

b. m(HCl) = m(CaCl2) × 2 × Mr(HCl) / Mr(CaCl2) = 325 g × 2 × 36.5 g/mol / 111 g/mol = 213.7 g

c. m(CO2) = m(CaCO3) × Mr(CO2) / Mr(CaCO3) = 750 g × 44 g/mol / 100 g/mol = 330 g